OpenStudy (studygirl606):

Not sure how to solve.... 3n/3=51/3

alones (alones):

http://mathpapa.com/

OpenStudy (jjcoolja12):

Let's solve your equation step-by-step. 3n 3 = 51 3 Step 1: Cross-multiply. 3n 3 = 51 3 3n*(3)=(51)*(3) 9n=153 Step 2: Divide both sides by 9. 9n 9 = 153 9 n=17 Answer: n=17

alones (alones):

@jjcoolja12 can you explain it clearly? that was fast looks like you dind't solve it all

rishavraj (rishavraj):

@jjcoolja12 dont give the whole solution in a go... let the user participate

OpenStudy (jjcoolja12):

guys if you've been asked to help multiple people at one you would go fast too

rishavraj (rishavraj):

@Studygirl606 u done with this question??

OpenStudy (studygirl606):

yeah

alones (alones):

@jjcoolja12 but you randomly post it, and liek the mod says, you suppose to mention the site you took it from, not copy and paste,

OpenStudy (jjcoolja12):

he is not a mod